1. A corporate bond has a 12 percent coupon, pays interest semiannually, and matures in 10 years at $1,000. If the investorsquestions 1-6 using financial calculator when possible
Answer #1

Solution :- 1

Face Value = $1000

As Semiannually So Coupon amount for 6 months = $1000 * 12% * 6/12 = $60

Time to Maturity = 10 years

As Semiannually so We take n = 10*2 = 20

Now Current Market Price of Bond = $60 * PVAF(7% , 20) + $1000 * PVF(7% , 20)

= ( $60 * 10.594 ) + ( $1000 * 0.258 )

= $635.64 + $258.419

= $894

As per HomeworkLib policy we need to answer only one question at once so please ask other as seperate one.